OpenStudy (anonymous):

domain of x+3/4-sqrt(x^2-9)??? explain answer please

OpenStudy (anonymous):

domain of a function is basically what values of x can be used in the function without computing an error or does not exist

OpenStudy (anonymous):

domain is x is an element of the Reals, you have no restrictions

OpenStudy (anonymous):

x cannot be 0 right because then it would be undefined

OpenStudy (anonymous):

yes correct, because you cannot take the sqrt of a negative unless your using complex numbers to simplify

OpenStudy (anonymous):

is this \[f(x)=\frac{x+3}{4}-\sqrt{x^2-9}\]?

OpenStudy (anonymous):

if so you have to make sure that \[x^2-9\geq0\] meaning \[x\leq -3\] or \[x\geq 3\]

OpenStudy (anonymous):

\[x+3/(4-\sqrt{x ^{2}-9})\]

OpenStudy (anonymous):

ok let then in that case not only must you have \[x\leq-3\] or \[x\geq 3\] you also have to make sure that \[\sqrt{x^2-9}\neq 4\] because you cannot divide by 0

OpenStudy (anonymous):

this means \[x^2-9\neq 16\] so \[x^2\neq 25\] so \[x\neq 5\] \[x\neq -5\]